Uranium-237 is a radioactive isotope with a half-life of 7*108 years. In a sample of initially 2*1015 atoms of 237U, how many are left after 2.1*109 years?

(In the "Answer" field, write the number of atoms in standard computer notation for powers of 10. For example, for 7*108 write "7E8".)


1
Expert's answer
2021-03-21T11:23:20-0400

The number of atoms that are left after "t = 2.1\times 10^{9}years" is given by the radioactive decay law:


"N(t) = N_02^{-t\/T}"

where "N_0 = 2\times 10^{15}" is the initial number of atoms, "T = 7\times 10^{8}years" is the half-life. Thus, obtain:


"N = 2\times 10^{15}\cdot 2^{-\left( \frac{2.1\times 10^{9}}{7\times 10^{8}} \right)} = 2.5\times 10^{14}"

Answer. 2.5E14.